I need to write a system of 2 equations in 2 unknowns for the problem.
At Gwen's garage sale all books were one price, and all magazines were another price. Harriet bouth 4 books and 3 magazines for $1.45, and June bought 2 books and 5 magazines for $1.25. What was the price of a book and what was the price of a magazine?

Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

write a system of 2 equations in 2 unknowns for the problem.
At Gwen's garage sale all books were one price, and all magazines were another price.
----------------------
Harriet bought 4 books and 3 magazines for $1.45,
June bought 2 books and 5 magazines for $1.25.
----------------------
What was the price of a book and what was the price of a magazine?
---
Equations:
4b + 3m = 145 cents
2b + 5m = 125 cents
======================
Cheers,
Stan H.

Answer by mananth(16949)

At Gwen's garage sale all books were one price, and all magazines were another price. Harriet bouth 4 books and 3 magazines for $1.45, and June bought 2 books and 5 magazines for $1.25. What was the price of a book and what was the price of a magazine?
let the price of a book be $x
AND the price of magazine be $y
four books will cost 4x & 3 magazines will cost 3y
4x+3y = 1.45
Similarly
2x+5y=1.25
2(2x+5y)-(4x+3y)= 2*1.25 - 1.45
4x+10y-4x-3y = 2.5 -1.45
7y= 1.05
y= 1.05/7
=0.15 Or 15 cents
4x+3y =1.45
plug the value of y
4x+3*0.15=1.45
4x+0.45 = 1.45-0.45
4x= 1
x= 0.25 OR 25 cents
Book cost 25 cents and magazine costs 15 cents